what is the 14th term in an arthmetic sequence in which the first term is 100 and the common diffence is -4?
Answer by Juhi(44)

we have a=100 and d=-4
nth term = a+(n-1)d
14th term = a+(14-1)d
= 100 + 13(-4)
= 100-52
= 48